OpenStudy (anonymous):

A circle is represented by the equation below: (x + 8)2 + (y − 3)2 = 100 Which statement is true? The circle is centered at (−8, 3) and has a radius of 20. The circle is centered at (8, −3) and has a diameter of 20. The circle is centered at (8, −3) and has a radius of 20. The circle is centered at (−8, 3) and has a diameter of 20.

OpenStudy (anonymous):

\(\large (x-h)^2+(y-k)^2=r^2 \), is the equation of a circle whose RADIUS is r, and center (h, k) according to the equation you gave, what's your radius? what's your center?

OpenStudy (anonymous):

I think that the answer is C because of one pair of (x + 8)2 being positive and (y − 3)2 being negative. @dpaInc

OpenStudy (anonymous):

no.... C says the radius is 20... it's not 20

OpenStudy (anonymous):

so would it be the diameter is 20 and the radius is 10? therefore making it B?

OpenStudy (anonymous):

the diameter is 20.... look at your equation carefully.... what's the center?

OpenStudy (anonymous):

I see the +8 and -3 so i thought it was those... do they switch because of the squaring?

OpenStudy (anonymous):

no, not because of the squaring... the equation have to written in standard form (with minus signs): \(\large (x+8)^2+(y-3)^2=100 \) \(\large (x--8)^2+(y-3)^2=10^2 \) \(\large (x-(\color{red} {-8}))^2+(y-(\color{red}3))^2=(\color{red}{10})^2 \) so that's how you get the center...

OpenStudy (anonymous):

So the center is (8,-3) correct? with a diameter of 20?

OpenStudy (anonymous):

it's in RED... center is (-8, 3) with a diameter of 20

OpenStudy (anonymous):

oh. i saw the double negative so i thought it became a positive

OpenStudy (anonymous):

just remember, if it's written with a plus sign, you need to write it with a minus like: \(\large (x+8)^2=(x--8)^2 \) so h=-8 in this case

OpenStudy (anonymous):

ok thank you :) i may have a few more questions and you explain them very well can i tag you in them when i get to them?

OpenStudy (anonymous):

yep... no problem...

OpenStudy (anonymous):

thx